Two small negatively charged spheres spaced 35.00 cm apart have equal amounts of net charge. If the magnitude of the force of repulsion between them is 2.430 x 1021 N, How many excess electrons must be present on each sphere? (For this calculation, use ke-8.988 x 109 Nm2/C2 for Coulombs constant and use qo 1.602 x 10-19C for the charge of an electron).
Answer #1

let the charge on each sphere be q

F = kq^2 / r^2

where q = charge and r is the separation between the charges

2.43*10^-21 = 9*10^9*q^2 / 0.35^2

q = -1.82*10^-7 C

excess electron in sphere = -1.82*10^-7 / (-1.6*10^-19) = 1.14*10^12